# Numerical instability of non-Hermitian Hamiltonian evolution

### v2

@@ VI Dynamics requires higher simulation precision than spectrum
In practice, diagonalizing a non-normal matrix ($H$) is computationally much faster than calculating non-Hermitian evolution ($e^{-\mathrm{i}Ht}|\psi_{0}\rangle/\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert$). This raises an important question: if we obtain a reliable spectrum with a precision of $P$ digits (i.e., numerical errors are smaller than a threshold $\epsilon$), does this guarantee that the non-Hermitian evolution computed with the same precision $P$ will also be reliable, with numerical errors of the same order? Our results suggest that the answer is no. For example, for the Hatano-Nelson model with $L=100$, $J=1.0$, $\gamma=0.8$: as shown in Fig. 3(c), with digits $P=30$, the diagonalization results are completely real with numerical errors $|E_{j}(P=30)-E^{(j)}_{\text{OBC}})|<10^{-6}$, $\forall j$. However, the numerical errors of non-Hermitian evolution are far larger than the $10^{-6}$ as depicted in Fig. 16, leading to qualitatively incorrect simulation results. As demonstrated in Fig. S5(a2), although the energy spectrum is reliable, the eigenvectors still deviate from the analytical results. Specifically, as we claimed before, the theoretical values of wave functions near the left edge are too small, beyond the scope of the digit $30$. These numerical errors in eigenvectors can serve as a signature of the incorrectness of the non-Hermitian evolution. In other words, the unreliability of eigenvectors means that non-Hermitian evolution is also unreliable. This correspondence is particularly clear in the Hatano-Nelson model, where the absolute value of the wave functions behave approximately as $r^{x}$, and the condition number scales as $r^{L-1}$. Given that the absolute value of the wave function around the right edge is $O(1)$, the absolute value of the wave function near the left boundary is about $O(1)/r^{L}$. For the non-Hermitian evolution, it is known the condition number gives a tight upper bound of the $\lVert e^{-\mathrm{i}Ht}\rVert$, implying that the minimum value of $1/\lVert e^{-\mathrm{i}Ht}\rVert$ scales as $r^{-(L-1)}$. Therefore, the incorrectness of eigenvectors indicates that the non-Hermitian evolution will also be unreliable. In summary, obtaining a reliable spectrum alone is insufficient to ensure accurate non-Hermitian evolution. The accuracy of the eigenvectors plays a more crucial role in the numerical stability of non-Hermitian evolution. In other words, numerically correct evolution results require higher precision than numerically correct spectrum results.
@@
Finally, while the models presented here are solvable and thus suitable for benchmarking, most non-Hermitian models, such as disordered, high-dimensional, and interacting systems, cannot be solved exactly. Consequently, it is generally challenging to assess the reliability of the spectrum and non-Hermitian evolution. However, our work underscores the importance of non-normality in determining the numerical stability of both the spectrum and non-Hermitian evolution. The non-normality is a universal property for non-Hermitian systems with NHSE, and it can be effectively described by the condition number. Therefore, given a non-Hermitian Hamiltonian, we can begin by diagonalizing it for small system sizes to extract the relationship between the condition number and the system size. Then, we can extrapolate the system size to larger values and monitor the numerical magnitude of the condition number. If the condition number exceeds about $O(10^{16})$, we need to improve the default double precision to higher precision digit $P$, making sure the condition number remains smaller than $O(10^{P})$. Especially for the non-Hermitian systems with NHSE, the condition number always exponentially grows with the system size, so it is essential to estimate the magnitude of the condition number before numerical calculations.
@@ VII Conclusions and Discussions
In this work, we investigate the numerical instability of non-Hermitian Hamiltonian with NHSE, a phenomenon often overlooked in practice [35]. Our results demonstrate that numerical errors can significantly impact the spectra, wave functions, and non-Hermitian evolution. Mathematically, this numerical instability is attributed to the large condition number of the non-Hermitian Hamiltonian with NHSE, which grows exponentially with the system size. Our work demonstrates that a reliable spectrum alone is not sufficient to ensure accurate non-Hermitian evolution. Instead, it is crucial to verify the reliability of the wave functions. Alternatively, it is paramount to estimate the condition number before numerical computations. To obtain reliable numerical results, we must ensure that the condition number is smaller than the $O(10^{P})$, in which $P$ is the precision digit used.
@@
Moreover, the skin effect has also been studied in the Markovian open quantum systems. According to quantum trajectory theory [55], the non-Hermitian evolution studied in this work can be understood as a special trajectory devoid of quantum jumps. With well-designed Lindblad operators, recent studies have explored the skin effect within the Lindbladian framework, incorporating all possible quantum trajectories. Interestingly, for such Lindbladians, the steady state is also approximately half side of the system being occupied (half-filled), and the entanglement entropy also follows an area law for $\gamma\neq 0$ [56, 57, 58, 59, 60, 61]. The dynamics of the Lindbladian with skin effect have also been simulated with the Gaussian-state method and the condition number of the Lindbladian also exponentially grows with the system size, so it is also interesting to investigate whether there exists similar numerical instability.
Our findings are not limited to one-dimensional free systems with NHSE; similar numerical instability is expected in disordered, higher-dimensional, interacting systems and Lindbladians with skin effect [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75]. Finally, this extreme sensitivity is not only relevant for numerical simulations but also for experiments, which are inevitably affected by defects, disorder, or noise. We predict that such imperfections in experimental setups may also lead to results that deviate significantly from the clean limit.
@@ Acknowledgements.
Appendix A Gaussian state simulation method and observables
@@
$S(L,L/2)=-\sum_{j=1}^{L/2}\left[\lambda_{j}\ \text{log}\lambda_{j}+(1-\lambda_{j})\ \text{log}(1-\lambda_{j})\right],$ (S4)
@@
Appendix B Direct observation of the numerical errors
Here, we directly observe the occurrence of numerical errors in non-Hermitian evolution, aiming to understand the origin of the spurious middle region. The Gaussian-state simulation method consists of two main steps: acting on the operator $e^{-\mathrm{i}H}$ (more strictly $e^{-\mathrm{i}H\delta t}$) and subsequent $QR$ decomposition. Notably, the incorrect smooth middle region persists regardless of the time step $\delta t$. Therefore, for simplicity, we choose the parameters $\gamma=0.8,\ L=60,\ \delta t=10$, and compare results obtained using double precision and $50$-digit precision. It is known that the $j$-th column of the $U$ matrix represents the single-particle wave function of the $j$-th particle. The Fig. S1 implies that the particles will tend to localize at the right side. More importantly, the steady-state $U$ matrix with double precision markedly deviates from the results with digit $50$. Specifically speaking, the correct results of digit $50$ show that the $j$-th particle has a high probability to localize around the site $L-j+1$, leading to a half-side occupied configuration. On the contrary, the $U$ matrix of double precision fails to converge properly, allowing particles to localize on the left side, thus inducing the smooth middle region. As stated in the Appendix. A, all the observables, including correlation functions, density distributions, currents, and entanglement entropy, are computed from $U$. Consequently, we argue that the double-precision numerical results reported in Ref. [35] are unreliable whenever the condition number for the chosen Hamiltonian parameters exceeds $10^{16}$.
Figure S2: The numerical diagonalized wave functions of the Hatano-Nelson with different system sizes. The figures plot all the wave functions. The blue line is the numerical result of double precision with $\gamma=-0.8$, while the red line is the numerical result of the digit 50 with $\gamma=0.8$ (performing the reflection operation). It is obvious that the double-precision results of $\gamma=-0.8$ are more accurate than $\gamma=0.8$ in Fig. 4.
@@ Appendix B Direct observation of the numerical errors
Appendix C Additional numerical results
@@
In the main text, we discuss the potential connection between the numerical errors in the wave functions and the errors in non-Hermitian evolution computed with double precision. Here, we present additional numerical data in Fig. S4 and Fig. S5 to further substantiate this observation, highlighting the correlation between the spatial extent of errors in the single-particle wave functions and in non-Hermitian evolution computed with the specified precision (digit $20$ and $30$).

### v3

@@ VI Dynamics requires higher simulation precision than spectrum
In practice, diagonalizing a non-normal matrix ($H$) is computationally much faster than calculating non-Hermitian evolution ($e^{-\mathrm{i}Ht}|\psi_{0}\rangle/\lVert e^{-\mathrm{i}Ht}|\psi_{0}\rangle\rVert$). This raises an important question: if we obtain a reliable spectrum with a precision of $P$ digits (i.e., numerical errors are smaller than a threshold $\epsilon$), does this guarantee that the non-Hermitian evolution computed with the same precision $P$ will also be reliable, with numerical errors of the same order? Our results suggest that the answer is no. For example, for the Hatano-Nelson model with $L=100$, $J=1.0$, $\gamma=0.8$: as shown in Fig. 3(c), with digits $P=30$, the diagonalization results are completely real with numerical errors $|E_{j}(P=30)-E^{(j)}_{\text{OBC}})|<10^{-6}$, $\forall j$. However, the numerical errors of non-Hermitian evolution are far larger than the $10^{-6}$ as depicted in Fig. 16, leading to qualitatively incorrect simulation results. As demonstrated in Fig. S5(a2), although the energy spectrum is reliable, the eigenvectors still deviate from the analytical results. Specifically, as we claimed before, the theoretical values of wave functions near the left edge are too small, beyond the scope of the digit $30$. These numerical errors in eigenvectors can serve as a signature of the incorrectness of the non-Hermitian evolution. In other words, the unreliability of eigenvectors means that non-Hermitian evolution is also unreliable. This correspondence is particularly clear in the Hatano-Nelson model, where the absolute values of the wave functions behave approximately as $r^{x}$, and the condition number scales as $r^{L-1}$. Given that the absolute value of the wave function around the right edge is $O(1)$, the absolute value of the wave function near the left boundary is about $O(1)/r^{L}$. For the non-Hermitian evolution, it is known the condition number gives a tight upper bound of the $\lVert e^{-\mathrm{i}Ht}\rVert$, implying that the minimum value of $1/\lVert e^{-\mathrm{i}Ht}\rVert$ scales as $r^{-(L-1)}$. Therefore, the incorrectness of eigenvectors indicates that the non-Hermitian evolution will also be unreliable. In summary, obtaining a reliable spectrum alone is insufficient to ensure accurate non-Hermitian evolution. The accuracy of the eigenvectors plays a more crucial role in the numerical stability of non-Hermitian evolution. In other words, numerically correct evolution results require higher precision than numerically correct spectrum results.
@@
Finally, while the models presented here are solvable and thus suitable for benchmarking, most non-Hermitian models, such as disordered, high-dimensional, and interacting systems, cannot be solved exactly. Consequently, it is generally challenging to assess the reliability of the spectrum and non-Hermitian evolution. However, our work underscores the importance of non-normality in determining the numerical stability of both the spectrum and non-Hermitian evolution. The strong non-normality is a universal property for non-Hermitian systems with NHSE, and it can be effectively described by the condition number. Therefore, given a non-Hermitian Hamiltonian, we can begin by diagonalizing it for small system sizes to extract the relationship between the condition number and the system size. Then, we can extrapolate the system size to larger values and monitor the numerical magnitude of the condition number. If the condition number is on the order of $O(10^{N})$ and exceeds the double-precision limit of approximately $O(10^{16})$, higher numerical precision is required. To ensure the reliability of the results, one may select a working precision of $P=N+20$ digits. This consideration is particularly important for non-Hermitian systems exhibiting the NHSE, where the condition number typically grows exponentially with system size. It is therefore essential to estimate its magnitude before numerical computations. Appendix E presents examples of disordered and interacting Hatano-Nelson models, both of which suffer from numerical instability. These cases highlight the generality of such instability in NHSE systems and underscore the importance of monitoring the condition number to ensure reliable numerical results.
@@ VII Conclusions and Discussions
In this work, we investigate the numerical instability of non-Hermitian Hamiltonians with NHSE, a phenomenon often overlooked in practice [36]. Our results demonstrate that numerical errors can significantly impact the spectra, wave functions, and non-Hermitian evolution. Mathematically, this numerical instability is attributed to the large condition number of the non-Hermitian Hamiltonian with NHSE, which grows exponentially with the system size. Our work demonstrates that a reliable spectrum alone is not sufficient to ensure accurate non-Hermitian evolution. Instead, it is crucial to verify the reliability of the wave functions. Therefore, it is paramount to estimate the condition number before numerical computations. Assuming the condition number is on the order of $O(10^{N})$, we can set the precision digit to be $P=N+20$, serving as a practical guideline for choosing precision in simulations involving highly non-normal matrices.
@@
Moreover, the skin effect has also been studied in the Markovian open quantum systems. According to quantum trajectory theory [56], the non-Hermitian evolution studied in this work can be understood as a special trajectory devoid of quantum jumps. With well-designed Lindblad operators, recent studies have explored the skin effect within the Lindbladian framework, incorporating all possible quantum trajectories. Remarkably, in such Lindbladian systems, the steady state remains approximately half-filled—i.e., localized on one side of the system—and the entanglement entropy obeys an area law for $\gamma\neq 0$ [57, 58, 59, 60, 61, 62]. The dynamics of Lindbladians exhibiting the skin effect have also been simulated using Gaussian-state methods. Moreover, the condition number of the Lindbladian grows exponentially with system size. This observation raises the intriguing question of whether similar numerical instabilities arise in such open quantum systems.
In this work, we present one-dimensional free, disordered, and interacting Hatano–Nelson models to illustrate the generality of numerical instability. These findings suggest that such instabilities are likely to be widespread in non-Hermitian and Lindbladian systems exhibiting the skin effect, regardless of the presence of disorder, dimensionality, or interactions [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76]. Finally, this extreme sensitivity is not only relevant for numerical simulations but also for experiments, which are inevitably affected by defects, disorder, or noise. We predict that such imperfections in experimental setups may also lead to results that deviate significantly from the clean limit.
@@ Acknowledgements.
Appendix A Numerical asymmetry induced by $QR$ algorithm
In the main text, we observed that double-precision numerical results for $\gamma=-0.8$ are significantly more accurate than those for $\gamma=0.8$, even though the corresponding Hamiltonians are the same after a transposition. To investigate the origin of this numerical asymmetry, we simulate the diagonalization procedure. The LAPACK routines for unsymmetric matrices follow three main steps: (1) Balancing: This step applies a similarity transformation: $\tilde{H}=D^{-1}P^{-1}HPD$, where $D$ is a diagonal scaling matrix and $P$ is a permutation matrix (if necessary). (2) Hessenberg Reduction: The balanced matrix $\tilde{H}$ is reduced to upper Hessenberg form $H^{\prime}$, $H^{\prime}=Q^{T}\tilde{H}Q$, in which $Q$ is an orthogonal matrix. The goal is to concentrate most of the mass on the diagonal and first sub-diagonal. (3) $QR$ algorithm: Eigenvalues are extracted from the Heisenberg matrix using the implicit double-shift $QR$ iteration. The goal is to find the real Schur form—a quasi-upper-triangular matrix from which real and complex eigenvalues are obtained from $1\times 1$ and $2\times 2$ blocks, respectively.
Numerical experiments show that step 3, the $QR$ algorithm, is the source of the numerical asymmetry. In MATLAB, steps 1 and 2 correspond to the functions “balance” and “hess”, respectively, while step 3. The step (3) can be executed using the function “schur”. We consider the Hatano-Nelson model with parameters $L=40$ and $\gamma=\pm 0.8$. Using MATLAB’s “schur” function to perform the Schur decomposition, we observe that for $\gamma=0.8$, the resulting quasi-upper-triangular matrix contains many $2\times 2$ blocks along the diagonal, indicating complex eigenvalues. In contrast, for $\gamma=-0.8$, the quasi-upper triangular matrix is composed of $1\times 1$ blocks, corresponding to real eigenvalues. This observation confirms that the numerical asymmetry originates from the Schur decomposition. To further monitor the occurrence of the numerical asymmetry, we simulate the $QR$ and $RQ$ iterations—key components of the Schur decomposition. After $4000$ iterations, both Hamiltonians are transformed into quasi-upper-triangular matrices. Consistently, we observe that $\gamma=0.8$ yields a spectrum dominated by $2\times 2$ blocks, whereas $\gamma=-0.8$ yields $1\times 1$ blocks, reproducing the asymmetry seen in the “schur” output.
A natural question arises: Why does the $QR$ algorithm exhibit greater numerical stability for $\gamma=-0.8$ than for $\gamma=0.8$?
To address this, we examine the standard implementation of the $QR$ algorithm via Householder reflections. This method iteratively eliminates sub-diagonal elements column by column using a sequence of orthogonal transformations. At each step $k$, a Householder matrix $T_{k}$ is constructed to zero out the sub-diagonal elements of column $k$ of a $L\times L$ matrix $A$. The Householder reflector is given by
$T_{k}=I-2\dfrac{v_{k}v_{k}^{T}}{v_{k}^{T}v_{k}},$ (S1)
where $v_{k}$ is chosen so that $T_{k}A$ has zeros below the diagonal in column $k$. The overall orthogonal matrix is the product $Q=T_{1}T_{2}T_{3}\cdots T_{L-1}$, and the transformed matrix becomes upper triangular: $R=T_{L-1}T_{L-2}\cdots T_{1}A$.
We implement this $QR$ decomposition using Householder reflections with parameters $L=40$ and $\gamma=\pm 0.8$. For clarity, we introduce the following notation. For $\gamma=0.8$, let: $H^{\text{pos}}=Q^{\text{pos}}_{0}R^{\text{pos}}_{0}$, $H_{1}^{\text{pos}}=R^{\text{pos}}_{1}Q^{\text{pos}}_{1}$, $H_{1}^{\text{pos}}=Q^{\text{pos}}_{1}R^{\text{pos}}_{1}$, $H_{2}^{\text{pos}}=R^{\text{pos}}_{1}Q^{\text{pos}}_{1}$, and so on. Similarly, for $\gamma=-0.8$, let: $H^{\text{neg}}=Q^{\text{neg}}_{0}R^{\text{neg}}_{0}$, $H_{1}^{\text{neg}}=R^{\text{neg}}_{1}Q^{\text{neg}}_{1}$, $H_{1}^{\text{neg}}=Q^{\text{neg}}_{1}R^{\text{neg}}_{1}$, $H_{2}^{\text{neg}}=R^{\text{neg}}_{1}Q^{\text{neg}}_{1}$, and so on. To assess numerical stability, we perform $QR$ and $RQ$ iterations using both double-precision and digit $50$. For $\gamma=-0.8$, the matrices $H_{j}^{\text{neg}}$ agree closely between the two precisions, indicating numerical stability. In contrast, for $\gamma=0.8$, significant deviations arise in $H_{j}^{\text{pos}}$ after only a few iterations, indicating numerical insufficiency of double precision. This discrepancy suggests that $H_{j}^{\text{neg}}$ is inherently more stable under Householder reflections than $H_{j}^{\text{pos}}$. Indeed, we observe a structural difference between them. The non-zero elements in the last few columns of $H_{j}^{\text{neg}}$ tend to be concentrated in the lower-right triangular region, whereas for $H_{j}^{\text{pos}}$, they concentrate in the upper-right region. This distinction affects the stability of the Householder transformation. In particular, while computing $T_{k}$ for $H_{j}^{\text{pos}}$, we can encounter extremely small denominators $v_{k}^{T}v_{k}$ (small than $O(10^{-16})$), leading to numerical instability. Such small denominators are not observed for $H_{j}^{\text{neg}}$, which explains its enhanced numerical robustness.
In summary, our numerical experiments demonstrate that the asymmetry in the numerical stability between $\gamma=\pm 0.8$ arises from structural differences in the matrices generated during $QR$ and $RQ$ iterations. The matrices $H_{j}^{\text{pos}}$ associated with $\gamma=0.8$ are more prone to instability due to the emergence of extremely small denominators in the Householder reflection process. In contrast, the structure of $H_{j}^{\text{neg}}$ for $\gamma=-0.8$ avoids this issue, resulting in greater numerical stability.
Appendix B Gaussian state simulation method and observables
@@
$S(L,L/2)=-\sum_{j=1}^{L/2}\left[\lambda_{j}\ \ln\lambda_{j}+(1-\lambda_{j})\ \ln(1-\lambda_{j})\right],$ (S5)
@@
Appendix C Direct observation of the numerical errors
Here, we directly observe the occurrence of numerical errors in non-Hermitian evolution, aiming to understand the origin of the spurious middle region. The Gaussian-state simulation method consists of two main steps: acting on the operator $e^{-\mathrm{i}H}$ (more strictly $e^{-\mathrm{i}H\delta t}$) and subsequent $QR$ decomposition. Notably, the incorrect smooth middle region persists regardless of the time step $\delta t$. Therefore, for simplicity, we choose the parameters $\gamma=0.8,\ L=60,\ \delta t=10$, and compare results obtained using double precision and $50$-digit precision. It is known that the $j$-th column of the $U$ matrix represents the single-particle wave function of the $j$-th particle. The Fig. S1 implies that the particles will tend to localize at the right side. More importantly, the steady-state $U$ matrix with double precision markedly deviates from the results with digit $50$. Specifically speaking, the correct results of digit $50$ show that the $j$-th particle has a high probability to localize around the site $L-j+1$, leading to a half-side occupied configuration. On the contrary, the $U$ matrix of double precision fails to converge properly, allowing particles to localize on the left side, thus inducing the smooth middle region. As stated in the Appendix. B, all the observables, including correlation functions, density distributions, currents, and entanglement entropy, are computed from $U$. Consequently, we argue that the double-precision numerical results reported in Ref. [36] are unreliable whenever the condition number for the chosen Hamiltonian parameters exceeds $10^{16}$.
Figure S2: The numerical diagonalized wave functions of the Hatano-Nelson with different system sizes. All wave functions are plotted: the blue lines represent double-precision numerical results for $\gamma=-0.8$, while the dashed green lines indicate the theoretical exponential localization predicted by the model, given by $|\psi(x)|=ce^{-x/\xi}$, where the localization length is $\xi=1/\ln\sqrt{(J+\gamma)/(J-\gamma)}$ and the prefactor $c$ depends on the system size. It should be noted that the dashed green lines are not the exact results of wave functions; they only approximately capture the exponential localization behaviors of the wave functions. It is evident that the double-precision results of $\gamma=-0.8$ are much more accurate than $\gamma=0.8$ in Fig. 4.
@@ Appendix C Direct observation of the numerical errors
Appendix D Additional numerical results
@@
In the main text, we discuss the potential connection between the numerical errors in the wave functions and the errors in non-Hermitian evolution computed with double precision. Here, we present additional numerical data in Fig. S4 and Fig. S5 to further substantiate this observation, highlighting the correlation between the spatial extent of errors in the single-particle wave functions and in non-Hermitian evolution computed with the specified precision (digits $20$ and $30$).
Appendix E Numerical instability of non-Hermitian disordered and interacting systems
We further investigate the disordered and interacting Hatano-Nelson models to demonstrate the generality of the observed numerical instability. The disordered Hatano-Nelson model is given by
$H_{\text{dis}}=\sum_{j}[(J+\gamma)c^{\dagger}_{j+1}c_{j}+(J-\gamma)c_{j}^{\dagger}c_{j+1}+\omega_{j}c_{j}^{\dagger}c_{j}],$ (S7)
where $w_{j}$ is a random on-site potential drawn from a uniform distribution $\omega_{j}\in[-W,W]$. Under open boundary conditions (OBC), the spectrum of this disordered Hatano-Nelson model remains real, as the Hamiltonian can be transformed into a Hermitian form via a similarity transformation $Q$, as introduced in Eq. (2). It is well known that wavefunctions of the disordered Hatano-Nelson model transition from skin localization to Anderson localization with the increase in disorder strength $W$ [63, 40]. As shown in Fig. S6, the numerical instability persists in the weak disorder regime, where the condition number grows exponentially with the system size. On the contrary, the strong disorder leads to Anderson localization, and the numerical instability is eliminated due to the small condition number [40]. For example, in Fig. S6(b1) and (b2), choosing parameters $W=0.5$, $L=50$, the condition number is much larger than $O(10^{16})$, so the density evolution displays the numerical instability as expected. Furthermore, we have also checked that the corresponding computed spectrum and wavefunctions are numerically unstable (not shown).
Figure S6: The disordered Hatano-Nelson model under OBC. $\gamma=0.9$, the data is averaged over 50 samples. (a) The condition number for different disorder strengths and system sizes. (b1) The density evolution with double precision. $W=0.5$, $L=50$. (b2) The density evolution with digit 40. $W=0.5$, $L=50$.
The interacting Hatano-Nelson model is defined as
$H_{\text{int}}=\sum_{j}[(J+\gamma)c_{j+1}^{\dagger}c_{j}+(J-\gamma)c^{\dagger}_{j}c_{j+1}+U_{\text{int}}n_{j}n_{j+1}],$ (S8)
in which $U_{\text{int}}$ denotes the nearest-neighbor interaction. The spectrum of the interacting Hatano-Nelson model remains real, since the model can also be mapped into a Hermitian Hamiltonian through a similarity transformation $Q_{\text{int}}=\text{exp}(\sum_{j}gjn_{j})$ with $e^{g}=r=\sqrt{(J+\gamma)/(J-\gamma)}$. This transformation yields
$Q_{\text{int}}^{-1}H_{\text{int}}Q_{\text{int}}=\sum_{j}\sqrt{J^{2}-\gamma^{2}}(c_{j}^{\dagger}c_{j+1}+\text{h.c.})+U_{\text{int}}n_{j}n_{j+1}.$ (S9)
As shown in Fig. S7(a), the condition number for the half-filled case scales exponentially with $L^{2}$. In particular, for the extreme non-reciprocity cases, i.e., $\gamma$ close to $1$, although the system size is limited for exact diagonalization, the condition number can still exceed $O(10^{16})$. Therefore, the default double precision becomes insufficient, resulting in inaccurate spectra and wavefunctions, as illustrated in Fig. S7(b). Moreover, as shown in Fig. S7(c), the norm $\lVert e^{-\mathrm{i}H_{\text{int}}t}\rVert$ can be much larger than $O(10^{16})$, so the non-Hermitian evolution will also be numerically incorrect with double precision.
In conclusion, numerical instability is a general feature of systems exhibiting the non-Hermitian skin effect (NHSE), regardless of the presence of disorder or interactions. The key underlying issue is the exponential scaling of the condition number with system size. Therefore, it is essential to estimate the condition number for small systems and extrapolate it to larger sizes to determine the required numerical precision for reliable computation.
Figure S7: The half-filled interacting Hatano-Nelson model under OBC. (a) The condition number for different non-reciprocity $\gamma$ and system sizes $L$. (b) The main plot depicts the absolute values of all the wavefunctions, which are rearranged from the largest to the smallest for better comparison. The $x$ axis of the main plot is the index of the Fock basis. The parameters are $\gamma=0.99$, $L=10$, particle number $N=5$. The size of the Hilbert space is $C_{L}^{N}=252$. [sxz: no mention on the inset] (c) The evolution of the norm $\lVert e^{-\mathrm{i}H_{\text{int}}t}\rVert$. $\gamma=0.999$, $L=10$, digit 50.